FIND THE FORMULA
A/B=C/D
D=4
C=6
B=8

Answer by unlockmath(1688)

Hello,
All we need to do is plug in the numbers with the letters like this:
A/B=C/D
D=4
C=6
B=8
4/8=6/D We need to solve what D is to cross multiply:
4D=48 Now divide both sides by 4 to give us:
D=12
Plug that into the original and you'll see it works.
RJ Toftness
www.math-unlock.com